compute the average velocity of the blood in the aorta of radius 1 cm if the flow rate is 5 liter/min
1
Expert's answer
2012-12-18T10:31:10-0500
compute the average velocity of the blood in the aorta of radius 1 cm if the flow rate is 5 liter/min

The area of the cross-section of the aorta of radius 1 cm is

S = π·(1[cm])² = π[cm²].

We'll take into account that

5[l] = 5·10³[cm³].

Therefore, the flow rate is

F = 5[l/min] = 5·10³[cm³/min]

and so the average velocity of the blood is

V = F/S = 5·10³[cm³/min] / (π[cm²]) ≈ 1591.55 [cm/min].